# 3Doodle: Compact Abstraction of Objects with 3D Strokes

### v1

@@ Appendix B Other Details
Dataset
We test the performance of 3Doodle in [Chair, Drums, Hotdog, Lego, Mic, Ship] scenes in the synthetic blender dataset from Neural Radiance Fields [29], and [Chair] scene from Invrender [51]. We additionally collected synthetic scenes to evaluate 3Doodle in various object shapes. We show the samples from our dataset in Figure 14. We slightly modify the publicly available 3D models from various sources below:
@@
Figure 14: Samples from our new dataset
Perceptual Study
@@
Implementation Details
In this section, we provide more details regarding rendering and optimizing our sketch parameters. To render the Bézier curve, we use the fixed width for the strokes in all of our experiments. However, naïve volume-rendering our $\sigma_{\text{contour}}$ leads non-uniform stroke widths in the $\epsilon_{i}$-$\alpha_{i}$ space. The small $\alpha_{i}$ and $\epsilon_{i}$ far from 1 makes the stroke thinner. Therefore, we set the width-related hyperparameters in $\mathcal{S}^{\text{3D}}_{\text{dep}}$ to be adaptive with respect to $\alpha_{i}$ and $\epsilon_{i}$:
$\gamma_{\text{adaptive}}=\begin{cases}\min\{f(\gamma_{\text{min}},\gamma_{\text{max}},\alpha_{i}),f(\gamma_{\text{min}},\gamma_{\text{max}},\epsilon_{i})\}&\epsilon_{i}<=1\\ \min\{f(\gamma_{\text{min}},\gamma_{\text{max}},\alpha_{i}),\gamma_{\text{max}}\}&\epsilon_{i}>1\end{cases},$ (25)
@@
$f(t_{\text{min}},t_{\text{max}},x)=t_{\text{min}}+(t_{\text{max}}-t_{\text{min}})\dfrac{x-0.1}{0.9}.$ (26)
@@
$b_{\text{adaptive}}=\begin{cases}\min\{f(b_{\text{min}},b_{\text{max}},\alpha_{i}),f(b_{\text{min}},b_{\text{max}},\epsilon_{i})\}&\epsilon_{i}<=1\\ \min\{f(b_{\text{min}},b_{\text{max}},\alpha_{i}),b_{\text{max}}\}&\epsilon_{i}>1\end{cases},$ (27)
$a_{\text{adaptive}}=\begin{cases}\max\{g(a_{\text{min}},a_{\text{max}},\alpha_{i}),\\ \qquad g(a_{\text{min}},a_{\text{max}},\epsilon_{i})\}&\alpha_{i},\epsilon_{i}>0.3\\ \max\{a_{\text{min}},a_{\text{max}},\alpha_{i})\}&\alpha_{i}>0.3,\epsilon_{i}<=0.3\\ \max\{g(a_{\text{min}},a_{\text{max}},\epsilon_{i}),a_{\text{min}}\}&\epsilon>0.3,\alpha<=0.3\\ a_{\text{min}}&\alpha_{i},\epsilon_{i}<=0.3,\end{cases},$ (28)
@@
$g(t_{\text{min}},t_{\text{max}},x)=t_{\text{max}}-(t_{\text{max}}-t_{\text{min}})\dfrac{x-0.1}{0.2}.$ (29)
When we optimize the scenes represented by both view-dependent sketches and view-independent sketches, we first optimize the superquadric parameters and then we optimize the Bézier curve parameters. Also, we do not apply robust loss for scenes that are represented by view-independent components solely.
Baselines
As we mentioned in the main manuscript, Kampelmühler and Pinz works for objects which are in the trained categories. Therefore, we evaluate them in the scenes that are in the car, teddy bear, hot air balloon, and chair categories. Also, NEF fails to reconstruct meaningful edge fields in our toycar and toyhorse scenes so we did not include the results in our quantitative results table.
Appendix C Code
We will release the code and dataset to evaluate our approach upon acceptance.
Appendix D Additional Results
We provide additional sketch generation results of our approach in Figure 15.
Figure 15: Additional results of our sketch

### v2

@@ Appendix B Other Details
B.1. Dataset
Figure 14. Samples from our new dataset
We test the performance of 3Doodle in [Chair, Drums, Hotdog, Lego, Mic, Ship] scenes in the synthetic blender dataset from Neural Radiance Fields (Mildenhall et al., 2021), and [Chair] scene from Invrender (Zhang et al., 2022b). We additionally collected synthetic scenes to evaluate 3Doodle in various object shapes. We show the samples from our dataset in Figure 14. We slightly modify the publicly available 3D models from various sources below:
@@ B.1. Dataset
B.2. Perceptual Study
@@
B.3. Implementation Details
In this section, we provide more details regarding rendering and optimizing our sketch parameters.
Rendering 3D Strokes
To render the Bézier curve, we use the fixed width for the strokes in all of our experiments. However, naïve volume-rendering our $\sigma_{\text{contour}}$ leads non-uniform stroke widths in the $\epsilon_{i}$-$\alpha_{i}$ space. The small $\alpha_{i}$ and $\epsilon_{i}$ far from 1 makes the stroke thinner. Therefore, we set the width-related hyperparameters in $\mathcal{S}^{\text{3D}}_{\text{dep}}$ to be adaptive with respect to $\alpha_{i}$ and $\epsilon_{i}$:
(17) $\gamma_{\text{adaptive}}=\begin{cases}\min\{f(\gamma_{\text{min}},\gamma_{\text{max}},\alpha_{i}),f(\gamma_{\text{min}},\gamma_{\text{max}},\epsilon_{i})\}&\epsilon_{i}<=1\\ \min\{f(\gamma_{\text{min}},\gamma_{\text{max}},\alpha_{i}),\gamma_{\text{max}}\}&\epsilon_{i}>1\end{cases},$
@@ B.3. Implementation Details
(18) $f(t_{\text{min}},t_{\text{max}},x)=t_{\text{min}}+(t_{\text{max}}-t_{\text{min}})\dfrac{x-0.1}{0.9}.$
@@
(19) $b_{\text{adaptive}}=\begin{cases}\min\{f(b_{\text{min}},b_{\text{max}},\alpha_{i}),f(b_{\text{min}},b_{\text{max}},\epsilon_{i})\}&\epsilon_{i}<=1\\ \min\{f(b_{\text{min}},b_{\text{max}},\alpha_{i}),b_{\text{max}}\}&\epsilon_{i}>1\end{cases},$
(20) $a_{\text{adaptive}}=\begin{cases}\max\{g(a_{\text{min}},a_{\text{max}},\alpha_{i}),\\ \qquad g(a_{\text{min}},a_{\text{max}},\epsilon_{i})\}&\alpha_{i},\epsilon_{i}>0.3\\ \max\{a_{\text{min}},a_{\text{max}},\alpha_{i})\}&\alpha_{i}>0.3,\epsilon_{i}<=0.3\\ \max\{g(a_{\text{min}},a_{\text{max}},\epsilon_{i}),a_{\text{min}}\}&\epsilon>0.3,\alpha<=0.3\\ a_{\text{min}}&\alpha_{i},\epsilon_{i}<=0.3,\end{cases},$
@@
(21) $g(t_{\text{min}},t_{\text{max}},x)=t_{\text{max}}-(t_{\text{max}}-t_{\text{min}})\dfrac{x-0.1}{0.2}.$
Optimization
When we optimize the scenes represented by both view-dependent sketches and view-independent sketches, we first optimize the superquadric parameters and then we optimize the Bézier curve parameters. Also, we do not apply robust loss for scenes that are represented by view-independent components solely. We initialize Bézier curves by sampling control points for the pre-defind number of stroke primitives. The control points for the real-world scenes and the teddy bear are randomly sampled from a bounding box. For all other objects, we use farthest point sampling on the SfM point cloud. The centers of superquadrics are also initialized by applying the farthest point sampling on the SfM points.
Line SfM-based Initialization
Figure 15. Reconstructed line maps for initialization.
In this subsection, we describe the details of our experiments of initialization method with line-based SfM. First, we obtain the 3D line maps from the known camera poses by triangulating the detected and matched line segments using LIMAP library (Liu et al., 2023). We show the reconstructed line maps in Fig. 15. The line-based SfM method can reconstruct the 3D line maps containing the structural details when the view-consistent edges are detectable (e.g., two chairs and lego). However, it struggles to reconstruct valid 3D line maps when the objects contain smooth contours (teddy bear and rocking pony), and consequently it leads to suboptimal results as demonstrated in Fig. 10 of the main manuscript. To initialize our sketch parameters with the linemaps, we first sample the line segments using the farthest sampling method based on the distance between lines. After lines are sampled, we use the centers of lines as the initial center of superquadrics. We also extract random samples along the lines to initialize the four control points of Bèzier curves.